# Learning to Solve Integer Linear Programs with Davis-Yin Splitting

### v3

@@ 4 DYS-Net
This approach works for simple sets $\mathcal{C}$ for which there exists an explicit form of $P_{\mathcal{C}}$, e.g. when $\mathcal{C}$ is the probability simplex Duchi et al. (2008); Condat (2016); Li et al. (2023). However, for general polytopes $\mathcal{C}$ no such form exists, thereby requiring a second iterative procedure, run at each iteration $k$, to compute $P_{\mathcal{C}}(x^{k})$. To avoid this computational expense, we draw on recent advances in the convex optimization literature [^3] See Pedregosa & Gidel (2018); Yurtsever et al. (2021) for similar applications of DYS, particularly Davis-Yin splitting or DYS Davis & Yin (2017). Specifically, we adapt the architecture incorporating DYS proposed in McKenzie et al. (2021). To this end, we rewrite $\mathcal{C}$ as
@@
If $\mathcal{C}_{1},\mathcal{C}_{2}$ are as in (14) and $A$ is full-rank then:
@@
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in (14), and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
@@
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
@@
for $F_{\Theta}(\cdot,d)$ into the formula for $T_{\Theta}(\cdot)$ given in (McKenzie et al., 2021, Theorem 3.2) and rearranging yields (16). Verifying the assumptions of (McKenzie et al., 2021, Theorem 3.3) are satisfied is straightforward. For example, $\nabla_{z}f_{\Theta}(z;\gamma,d)$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem Baillon & Haddad (1977); Bauschke & Combettes (2009).
The forward pass of DYS-net consists of iterating (15) until a suitable stopping criterion is met, which for simplicity we assume to be a maximum number of iterations.
@@
From Theorem 2 we deduce the following fixed point condition:
@@
As discussed in McKenzie et al. (2021), instead of backpropagating through every iterate of the forward pass, we may derive a formula for $dx_{\Theta}/d\Theta$ by appealing to the implicit function theorem and differentiating both sides of (17):
@@
We notice two immediate problems not addressed by McKenzie et al. (2021): (i) $T_{\Theta}$ is not everywhere differentiable with respect to $z$, as $P_{\mathcal{C}_{2}}$ is not; (ii) if $T_{\Theta}$ were a contraction (i.e. Lipschitz in $z$ with constant less than $1$), then $\mathcal{J}_{\Theta}$ would be invertible. However, this is not necessarily the case. Thus, it is not clear a priori that (18) can be solved for $\mbox{d}{{z}_{\Theta}}/\mbox{d}{\Theta}$. Our key result (Theorem 6 below) is to provide reasonable conditions under which $\mathcal{J}_{\Theta}(z_{\Theta})$ is invertible.
@@
$\frac{\mathrm{d}\ell}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}{x}_{\Theta}}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\left(\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}\right)=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\mathcal{J}_{\Theta}^{-1}\frac{\partial T_{\Theta}}{\partial\Theta}$
This approach requires solving a linear system with $\mathcal{J}_{\Theta}$ which becomes particularly expensive when $n$ is large. Instead, we use Jacobian-free Backpropagation (JFB) Fung et al. (2022), also known as one-step differentiation Bolte et al. (2024), in which the Jacobian $\mathcal{J}_{\Theta}$ is replaced with the identity matrix. This leads to an approximation of the true gradient $\mbox{d}{\ell}/\mbox{d}{\Theta}$ using
$p_{\Theta}(x_{\Theta})=\left[\frac{\partial\ell}{\partial x}\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\frac{\partial T_{\Theta}}{\partial\Theta}\right]_{(x,z)=(x_{\Theta},z_{\Theta})}.$ (19)
We show (19) is a valid descent direction by resolving the two problems highlighted above. We begin by rigorously deriving a formula for $\partial{T_{\Theta}}/\partial{z}$. Recall the following generalization of the Jacobian to non-smooth operators due to Clarke (1983).
@@
For any locally Lipshitz $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ let $D_{F}$ denote the set upon which $F$ is differentiable. The Clarke Jacobian of $F$ is the set-valued function defined as
@@
This aligns with the default rule for assigning a sub-gradient to $\operatorname{ReLU}$ used in the popular machine learning libraries TensorFlowAbadi et al. (2016), PyTorch Paszke et al. (2019) and JAX Bradbury et al. (2018), and has been observed to yield networks which are more stable to train than other choices Bertoin et al. (2021).
@@
2:  Initialize: Attach neural network $w_{\Theta}(\cdot)$. Compute SVD of $A$ for $P_{\mathcal{C}_{1}}$ formula
3:  Forward Pass: Initialize $z^{0}$. Given $w_{\Theta}(d)$ compute $z^{1},\ldots,z^{K}$ using (15). Return $x^{K}\triangleq P_{\mathcal{C}_{1}}(z^{K})\approx{x}_{\Theta}$.
4:  Backward Pass: Compute $p_{\Theta}(x^{K})\approx p_{\Theta}({x}_{\Theta})$ using (19) and return.
@@
Let ${x}_{\Theta}$ denote the solution to (8). Let $\mathcal{A}({x}_{\Theta})\subseteq\{1,\ldots,n\}$ denote the set of active positivity constraints:
@@
Theorem 6 also proves that $\mathcal{J}_{\Theta}(z_{\Theta})$ in (18) is invertible, thus also justifying the use of Jacobian-based backprop as well as numerous other gradient approximation techniques Liao et al. (2018); Geng et al. (2021).
Finally, we note that in practice $p_{\Theta}(x^{K})$ is used instead of $p_{\Theta}(x_{\Theta})$. (Fung et al., 2022, Corollary 0.1) guarantees that $p_{\Theta}(x^{K})$ will still be a descent direction for $K$ sufficiently large, see also (Bolte et al., 2024, Cor. 1).
@@
DYS-net is implemented as an abstract PyTorch Paszke et al. (2019) layer in the provided code. To instantiate it, a user only need specify $A$ and $b$ (see (3)) and choose an architecture for $w_{\Theta}$. At test time, it can be beneficial to solve the underlying combinatorial problem (2) exactly using a specialized algorithm, e.g. commercial integer programming software for the knapsack prediction problem. This can easily be done using DYS-net by instantiating the test-time-forward abstract method.
@@ 5.1 Experiments using PyEPO
First, we verify the efficiency of DYS-net by performing benchmarking experiments within the PyEPO Tang & Khalil (2022) framework. We compare DYS-net against three other methods: Perturbed Optimization Berthet et al. (2020)PertOpt-net; an approach using Blackbox Backpropagation Vlastelica et al. (2019)BB-net; and an approach using cvxpylayers Agrawal et al. (2019a)CVX-net.
@@
We use the PyEPO implementations of Perturbed Optimization, respectively Blackbox Backpropagation, as the core of PertOpt-net, respectively BB-net. We use the same architecture for $w_{\Theta}(d)$ for all models [^4] The architecture we use for shortest path prediction problems differs from that used for knapsack prediction problems. See the appendix for details. and use an appropriate combinatorial solver at test time, via the PyEPO interface to Gurobi. Thus, the methods differ only in training procedure. The precise architectures used for $w_{\Theta}(d)$ for each problem are described in Appendix C, and are also viewable in the accompanying code repository [^5] See code at https://github.com/mines-opt-ml/fpo-dys .Interestingly, we observed that adding drop-out during training to the output layer proved beneficial for the knapsack prediction problem. Without this, we find that $w_{\Theta}$ tends to output a sparse approximation to $w$ supported on a feasible set of items, and so does not generalize well.
@@
We train for a maximum of 100 epochs or 30 minutes, whichever comes first. We use a validation set for model selection as we observe that, for all methods, the best loss is seldom achieved at the final iteration. Exact hyperparameters are discussed in the appendix. For each problem size, we perform three repeats with different randomly generated data. Results are displayed in Figure 2.
@@
Given test data $\{(d_{i},w(d_{i}),x^{\star}(d_{i}))\}_{i=1}^{N}$ and a fully trained model $w_{\Theta}(d)$ we define the regret as:
@@ 5.1.1 Shortest Path Prediction
where $E$ is the vertex-edge adjacency matrix, $b$ encodes the initial and terminal vertices, and $w(d)\in\mathbb{R}^{|\mathcal{E}|}$ is a vector encoding ($d$-dependent) edge lengths; see Appendix for further details. Here, $x^{\star}$ will encode the edges included in the optimal path. In this experiment we focus on the case where $\mathcal{G}$ is the $k\times k$ grid graph. We use the PyEPO Tang & Khalil (2022) benchmarking software to generate training datasets of the form $\{(d_{i},x^{\star}_{i}\approx x^{\star}(d_{i}))\}_{i=1}^{N}$ where $N=1000$ for each $k\in\{5,10,15,20,25,30\}$. The $d_{i}$ are sampled from the five-dimensional multivariate Gaussian distribution with mean 00 and identity covariance. Note that the number of variables in (26) scales like $k^{2}$, not $k$.
@@ 5.1.3 Results
As is clear from Figure 2 DYS-net trains the fastest among the four methods, and achieves the best or near-best normalized regret in all cases.
@@ 5.2 Large-scale shortest path prediction
Our shortest path prediction experiment described in Section 5.1.1 is bottlenecked by the fact that the “academic” license of Gurobi does not allow for shortest path prediction problems on grids larger than $30\times 30$. To further explore the limits of DYS-net, we redo this experiment using a custom pyTorch implementation of Dijkstra’s algorithm666adapted from Tensorflow code available at https://github.com/google-research/google-research/blob/master/perturbations/experiments/shortest˙path.py as our base solver for (26).
@@
We generate datasets $\mathcal{D}=\{(d_{i},x^{\star}(d_{i})\}_{i=1}^{1,000}$ for $k$-by-$k$ grids where $k\in\{5,10,20,30,50,100\}$ and the $d_{i}$ are sampled uniformly at random from $[0,1]^{5}$, the true edge weights are computed as $w(d)=Wd$ for fixed $W\in\mathbb{R}^{|\mathcal{E}|\times 5}$, and $x^{\star}(d)$ is computed given $w(d)$ using the aforementioned pyTorch implementation of Dijkstra’s algorithm. Further details are presented in Appendix C.
@@
grid size number of variables network size
@@
Figure 4: Results for the shortest path prediction problem. a) Test MSE loss (left), b) training time in minutes (middle), and c) regret values (right) vs. gridsize for DYS-Net (proposed) and approaches using cvxpylayers Agrawal et al. (2019b) labeled CVX; Perturbed Optimization Berthet et al. (2020) labeled PertOpt; and Blackbox Backpropagation Vlastelica et al. (2019), labeled BB. Note CVX is unable to load or run problems with gridsize over 30. Dimensions of the variables can be found in Table 1.
@@
We test the same four approaches as in Sections 5.1.1 and 5.1.2, but unlike in these sections we do not use the PyEPO implementations of PertOpt-net or BB-net. We cannot, as the PyEPO implementations call Gurobi to solve (26) which, as mentioned above, cannot handle grids larger than $30\times 30$. Instead, we use custom implementations based on the code accompanying the works introducing PertOpt Berthet et al. (2020) and BB Vlastelica et al. (2019). We use the same architecture for $w_{\Theta}(d)$ for DYS-net, PertOpt-net, and Cvx-net; a two layer fully connected neural network with leaky ReLU activation functions. For DYS-net and Cvx-net we do not modify the forward pass at test time. For BB-net we use a larger network by making the latent dimension 20-times larger than that of the first three as we found this more effective. Network sizes can be seen in Table 1.
We tuned the hyperparameters for each approach to the best of our ability on the smallest problem (5-by-5 grid) and then used these hyperparameter values for all other graph sizes. See Figure 3 for the results of this training run. We train all approaches for 100 epochs total on each problem using the $\ell_{2}$ loss (5).

### v4

@@ 4 DYS-Net
This approach works for simple sets $\mathcal{C}$ for which there exists an explicit form of $P_{\mathcal{C}}$, e.g. when $\mathcal{C}$ is the probability simplex (Duchi et al., 2008; Condat, 2016; Li et al., 2023). However, for general polytopes $\mathcal{C}$ no such form exists, thereby requiring a second iterative procedure to compute $P_{\mathcal{C}}(x^{k})$. This projection must be computed at every iteration of every forward pass for every sample of every epoch, and this dominates the computational cost, see McKenzie et al. (2021, Appendix 4).
To avoid this expense, we draw on recent advances in the convex optimization literature, particularly Davis-Yin splitting or DYS (Davis & Yin, 2017). DYS is a three operator splitting technique, and hence is able to decouple the polytope constraints into relatively simple constraints Davis & Yin (2017). This technique has been used elsewhere, e.g. Pedregosa & Gidel (2018); Yurtsever et al. (2021). Specifically, we adapt the architecture incorporating DYS proposed in McKenzie et al. (2021). To this end, we rewrite $\mathcal{C}$ as
@@
If $\mathcal{C}_{1},\mathcal{C}_{2}$ are as in equation 14 and $A$ is full-rank then:
@@
We remark that the SVD of $A$ is computed offline, and is a once-off cost. Alternatively, one could compute $P_{\mathcal{C}_{1}}$ by using an iterative method (e.g. GMRES) to find the (least squares) solution to $Av=Az-b$, whence $P_{\mathcal{C}_{1}}(z)=z-v$, but this would need to be done at each iteration of each forward pass for every sample of every epoch. Thus, it is unclear whether this procides computational savings over the once-off computation of the SVD. However, this could prove useful when $A$ is only provided implicitly, i.e. via access to a matrix-vector product oracle.
@@
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in equation 14, and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
@@
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then the sequence $\{x^{k}\}$ converges to $x_{\Theta}$ in equation 8 and $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
@@
for $F_{\Theta}(\cdot,d)$ into the formula for $T_{\Theta}(\cdot)$ given in McKenzie et al. (2021, Theorem 3.2) and rearranging yields equation 16. Verifying the assumptions of McKenzie et al. (2021, Theorem 3.3) are satisfied is straightforward. For example, $\nabla_{z}f_{\Theta}(z;\gamma,d)$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem (Baillon & Haddad, 1977; Bauschke & Combettes, 2009).
The forward pass of DYS-net consists of iterating equation 15 until a suitable stopping criterion is met. For simplicity, in presenting Algorithm 1 we assume this to be a maximum number of iterations is reached. One may also consider a stopping criterion of $\|z_{k+1}-z_{k}\|\leq\mathrm{tol}$.
@@
From Theorem 2, we deduce the following fixed point condition:
@@
As discussed in McKenzie et al. (2021), instead of backpropagating through every iterate of the forward pass, we may derive a formula for $dx_{\Theta}/d\Theta$ by appealing to the implicit function theorem and differentiating both sides of equation 17:
@@
We notice two immediate problems not addressed by McKenzie et al. (2021): (i) $T_{\Theta}$ is not everywhere differentiable with respect to $z$, as $P_{\mathcal{C}_{2}}$ is not; (ii) if $T_{\Theta}$ were a contraction (i.e. Lipschitz in $z$ with constant less than $1$), then $\mathcal{J}_{\Theta}$ would be invertible. However, this is not necessarily the case. Thus, it is not clear a priori that equation 18 can be solved for $\mbox{d}{{z}_{\Theta}}/\mbox{d}{\Theta}$. Our key result (Theorem 6 below) is to provide reasonable conditions under which $\mathcal{J}_{\Theta}(z_{\Theta})$ is invertible.
@@
$\frac{\mathrm{d}\ell}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}{x}_{\Theta}}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\left(\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}\right)=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\mathcal{J}_{\Theta}^{-1}\frac{\partial T_{\Theta}}{\partial\Theta}$
This approach requires solving a linear system with $\mathcal{J}_{\Theta}$ which becomes particularly expensive when $n$ is large. Instead, we use Jacobian-free Backpropagation (JFB) (Fung et al., 2022), also known as one-step differentiation (Bolte et al., 2024), in which the Jacobian $\mathcal{J}_{\Theta}$ is replaced with the identity matrix. This leads to an approximation of the true gradient $\mbox{d}{\ell}/\mbox{d}{\Theta}$ using
$p_{\Theta}(x_{\Theta})=\left[\frac{\partial\ell}{\partial x}\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\frac{\partial T_{\Theta}}{\partial\Theta}\right]_{(x,z)=(x_{\Theta},z_{\Theta})}.$ (19)
This update can be seen as a zeroth order Neumann expansion of $\mathcal{J}_{\Theta}^{-1}$ Fung et al. (2022); Bolte et al. (2024). We show equation 19 is a valid descent direction by resolving the two problems highlighted above. We begin by rigorously deriving a formula for $\partial{T_{\Theta}}/\partial{z}$. Recall the following generalization of the Jacobian to non-smooth operators due to Clarke (1983).
@@
For any locally Lipschitz $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ let $D_{F}$ denote the set upon which $F$ is differentiable. The Clarke Jacobian of $F$ is the set-valued function defined as
@@
This aligns with the default rule for assigning a sub-gradient to $\operatorname{ReLU}$ used in the popular machine learning libraries TensorFlow(Abadi et al., 2016), PyTorch (Paszke et al., 2019) and JAX (Bradbury et al., 2018), and has been observed to yield networks which are more stable to train than other choices (Bertoin et al., 2021).
@@
2:  Initialize: Attach neural network $w_{\Theta}(\cdot)$. Compute SVD of $A$ for $P_{\mathcal{C}_{2}}$ formula
3:  Forward Pass: Initialize $z^{0}$. Given $w_{\Theta}(d)$ compute $z^{1},\ldots,z^{K}$ using equation 15. Return $x^{K}\triangleq P_{\mathcal{C}_{1}}(z^{K})\approx{x}_{\Theta}$.
4:  Backward Pass: Compute $p_{\Theta}(x^{K})\approx p_{\Theta}({x}_{\Theta})$ using equation 19 and return.
@@
Let ${x}_{\Theta}$ denote the solution to equation 8. Let $\mathcal{A}({x}_{\Theta})\subseteq\{1,\ldots,n\}$ denote the set of active positivity constraints:
@@
Theorem 6 also proves that $\mathcal{J}_{\Theta}(z_{\Theta})$ in equation 18 is invertible, thus also justifying the use of Jacobian-based backprop as well as numerous other gradient approximation techniques (Liao et al., 2018; Geng et al., 2021).
Finally, we note that in practice $p_{\Theta}(x^{K})$ is used instead of $p_{\Theta}(x_{\Theta})$. Fung et al. (2022, Corollary 0.1) guarantees that $p_{\Theta}(x^{K})$ will still be a descent direction for $K$ sufficiently large, see also Bolte et al. (2024, Cor. 1).
Synergy between forward and backward pass
We highlight that the pseudogradient computed in the backward pass (i.e., $p_{\Theta}$) does not require any Lagrange multipliers, in contrast with the gradient computed using equation 12. This is why we may use Davis-Yin splitting, as opposed to a primal-dual interior point method, on the forward pass. This allows efficiency gains on the forward pass, as Davis-Yin splitting scales favorably with the problem dimension.
@@
DYS-net is implemented as an abstract PyTorch (Paszke et al., 2019) layer in the provided code. To instantiate it, a user only need specify $A$ and $b$ (see equation 3) and choose an architecture for $w_{\Theta}$. At test time, it can be beneficial to solve the underlying combinatorial problem equation 2 exactly using a specialized algorithm, e.g. commercial integer programming software for the knapsack prediction problem. This can easily be done using DYS-net by instantiating the test-time-forward abstract method.
@@ 5 Numerical Experiments
We demonstrate the effectiveness of DYS-Net through four experiments. In all our experiments with DYS-net we used $\alpha=0.05$ and $K=1,000$, with an early stopping condition of $|z_{k+1}-z_{k}|_{2}\leq\mathrm{tol}$ with $\mathrm{tol}=0.01$. We compare DYS-net against three other methods: Perturbed Optimization (Berthet et al., 2020)PertOpt-net; an approach using Blackbox Backpropagation (Vlastelica et al., 2019)BB-net; and an approach using cvxpylayers (Agrawal et al., 2019a)CVX-net. All code needed to reproduce our experiments is available at https://github.com/mines-opt-ml/fpo-dys.
@@ 5.1 Experiments using PyEPO
First, we verify the efficiency of DYS-net by performing benchmarking experiments within the PyEPO (Tang & Khalil, 2022) framework.
@@
We use the PyEPO implementations of Perturbed Optimization, respectively Blackbox Backpropagation, as the core of PertOpt-net, respectively BB-net. We use the same architecture for $w_{\Theta}(d)$ for all models [^3] The architecture we use for shortest path prediction problems differs from that used for knapsack prediction problems. See the appendix for details. and use an appropriate combinatorial solver at test time, via the PyEPO interface to Gurobi. Thus, the methods differ only in training procedure. The precise architectures used for $w_{\Theta}(d)$ for each problem are described in Appendix C, and are also viewable in the accompanying code repository. Interestingly, we observed that adding drop-out during training to the output layer proved beneficial for the knapsack prediction problem. Without this, we find that $w_{\Theta}$ tends to output a sparse approximation to $w$ supported on a feasible set of items, and so does not generalize well.
@@
We train for a maximum of 100 epochs or 30 minutes, whichever comes first. We use a validation set for model selection as we observe that, for all methods, the best loss is seldom achieved at the final iteration. Exact hyperparameters are discussed in Appendix C. For each problem size, we perform three repeats with different randomly generated data. Results are displayed in Figure 2.
@@
Given test data $\{(d_{i},w(d_{i}),x^{\star}(d_{i}))\}_{i=1}^{N}$ and a fully trained model $w_{\Theta}(d)$ we define the regret:
@@ 5.1.1 Shortest Path Prediction
where $E$ is the vertex-edge adjacency matrix, $b$ encodes the initial and terminal vertices, and $w(d)\in\mathbb{R}^{|\mathcal{E}|}$ is a vector encoding ($d$-dependent) edge lengths. Here, $x^{\star}$ will encode the edges included in the optimal path. In this experiment we focus on the case where $\mathcal{G}$ is the $k\times k$ grid graph. We use the PyEPO Tang & Khalil (2022) benchmarking software to generate training datasets of the form $\{(d_{i},x^{\star}_{i}\approx x^{\star}(d_{i}))\}_{i=1}^{N}$ where $N=1000$ for each $k\in\{5,10,15,20,25,30\}$. The $d_{i}$ are sampled from the five-dimensional multivariate Gaussian distribution with mean 00 and identity covariance. Note that the number of variables in equation 26 scales like $k^{2}$, not $k$.
@@ 5.1.3 Results
As is clear from Figure 2 DYS-net trains the fastest among the four methods, and achieves the best or near-best normalized regret in all cases. We attribute the accuracy to the fact that it is not a zeroth order method. Moreover, avoiding computation and backpropagation through the projection $P_{\mathcal{C}}$ allows DYS to have fast runtimes.
@@ 5.2 Large-scale shortest path prediction
Our shortest path prediction experiment described in Section 5.1.1 is bottlenecked by the fact that the “academic” license of Gurobi does not allow for shortest path prediction problems on grids larger than $30\times 30$. To further explore the limits of DYS-net, we redo this experiment using a custom pyTorch implementation of Dijkstra’s algorithm444adapted from Tensorflow code available at https://github.com/google-research/google-research/blob/master/perturbations/experiments/shortest_path.py as our base solver for equation 26.
@@
We generate datasets $\mathcal{D}=\{(d_{i},x^{\star}(d_{i})\}_{i=1}^{1,000}$ for $k$-by-$k$ grids where $k\in\{5,10,20,30,50,100\}$ and the $d_{i}$ are sampled uniformly at random from $[0,1]^{5}$, the true edge weights are computed as $w(d)=Wd$ for fixed $W\in\mathbb{R}^{|\mathcal{E}|\times 5}$, and $x^{\star}(d)$ is computed given $w(d)$ using the aforementioned pyTorch implementation of Dijkstra’s algorithm. Thus, all entries of $w(d)$ are non-negative, and are positive with overwhelming probability.
@@
grid size number of variables neural network size
@@
Figure 4: Results for the shortest path prediction problem. a) Test MSE loss (left), b) training time in minutes (middle), and c) regret values (right) vs. gridsize for DYS-Net (proposed) and approaches using cvxpylayers (Agrawal et al., 2019b) labeled CVX; Perturbed Optimization (Berthet et al., 2020) labeled PertOpt; and Blackbox Backpropagation (Vlastelica et al., 2019), labeled BB. Note CVX is unable to load or run problems with gridsize over 30. Dimensions of the variables can be found in Table 1.
@@
We test the same four approaches as in Sections 5.1.1 and 5.1.2, but unlike in these sections we do not use the PyEPO implementations of PertOpt-net or BB-net. We cannot, as the PyEPO implementations call Gurobi to solve equation 26 which, as mentioned above, cannot handle grids larger than $30\times 30$. Instead, we use custom implementations based on the code accompanying the works introducing PertOpt (Berthet et al., 2020) and BB (Vlastelica et al., 2019). We use the same architecture for $w_{\Theta}(d)$ for DYS-net, PertOpt-net, and Cvx-net; a two layer fully connected neural network with leaky ReLU activation functions. For DYS-net and Cvx-net we do not modify the forward pass at test time. For BB-net we use a larger network by making the latent dimension 20-times larger than that of the first three as we found this more effective. Network sizes can be seen in Table 1.
We tuned the hyperparameters for each approach to the best of our ability on the smallest problem (5-by-5 grid) and then used these hyperparameter values for all other graph sizes. See Figure 3 for the results of this training run. We train all approaches for 100 epochs total on each problem using the $\ell_{2}$ loss equation 5.